OpenStudy (anonymous):

In a 3x3 matrix, how do I calculate the eigenvalues and eigenvectors, given that Ax=λx, λ=4? How do I find a, b or c? Thank?

OpenStudy (anonymous):

and what is a, b and c?

OpenStudy (anonymous):

To find eigen values: \[ \det(A-\lambda I)=0 \]Solve for \(\lambda\).

OpenStudy (anonymous):

Basically: 1) subtract \(\lambda\) from the diagonals 2) find the determinate, giving you a polynomial 3) find the roots of that polynomial

OpenStudy (anonymous):

Thx, wio. Let me try that.

OpenStudy (anonymous):

If you have repeated or imaginary eigen values, it gets really hairy. Hopefully they're all unique and real.

OpenStudy (anonymous):

Okay, I have -2, -1, -2 on the diagonal. Do i subtract λ diagonally and proceed or?

OpenStudy (anonymous):

In that case the eigen vector \(\mathbf v_i\) corresponding to the eigen value \(\lambda_i\) can be found by solving: \[ (A-\lambda_i I)\mathbf v_i = \mathbf 0 \]

OpenStudy (anonymous):

I think? Actually I'm pretty sure it's just finding the null space of: \[ (A-\lambda_i) \]

OpenStudy (anonymous):

Which may be the same thing as that equation above...

OpenStudy (anonymous):

\[ A-\lambda_i I \]

OpenStudy (anonymous):

I'm gettting confused here. Did you see the pic I attached?

OpenStudy (anonymous):

Yes, why?

OpenStudy (anonymous):

I understand I need to subtract λ diagonally, but why is it given?

OpenStudy (anonymous):

I don't know. It isn't even an Eigen value for the picture you uploaded.

OpenStudy (anonymous):

So, are you suggesting the question was wrong?

OpenStudy (anonymous):

That is a possibility.

OpenStudy (anonymous):

I think we might be interpreting it wrong, or you could have written it down wrong.

OpenStudy (anonymous):

Okay. Here is the full thing.

OpenStudy (anonymous):

Hello wio. Still there?

ganeshie8 (ganeshie8):

question looks fine to me, to make ur life simple, they gave u one eigen value. you're supposed to solve other two roots

ganeshie8 (ganeshie8):

@d3Xter you able to find the eigenvectors ?

OpenStudy (anonymous):

Well, using \[ \lambda I _{3}-A.\] I got \[\lambda ^{3}-7\lambda ^{2}+14\lambda-8\]

ganeshie8 (ganeshie8):

ok, \(\lambda = 2 \) satisfies that equation, so it is clar the thw question is asking us to find the eigenvector corresponding to \(\lambda = 2 \)

ganeshie8 (ganeshie8):

you knw how to find the corresponding eigenvector right ?

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

How do I go about that?

ganeshie8 (ganeshie8):

\[ \left( \begin{array}{ccc} 2-\lambda & 2 & -2 \\ 1 & 3-\lambda & 1 \\ 1 & 2 & 2-\lambda \end{array} \right) \left( \begin{array}{c} x \\ y \\ z \end{array} \right) = 0 \]

ganeshie8 (ganeshie8):

put \(-\lambda = 2\) and solve above

ganeshie8 (ganeshie8):

\[ \left( \begin{array}{ccc} 2-2 & 2 & -2 \\ 1 & 3-2 & 1 \\ 1 & 2 & 2-2 \end{array} \right) \left( \begin{array}{c} x \\ y \\ z \end{array} \right) = 0 \] \[ \left( \begin{array}{ccc} 0 & 2 & -2 \\ 1 & 1 & 1 \\ 1 & 2 & 0 \end{array} \right) \left( \begin{array}{c} x \\ y \\ z \end{array} \right) = 0 \]

ganeshie8 (ganeshie8):

reduce the matrix

ganeshie8 (ganeshie8):

\[ \left( \begin{array}{ccc} 0 & 2 & -2 & 0 \\ 1 & 1 & 1 & 0\\ 1 & 2 & 0 & 0 \end{array} \right) \] \[ \left( \begin{array}{ccc} 1 & 1 & 1 & 0\\ 0 & 2 & -2 & 0 \\ 1 & 2 & 0 & 0 \end{array} \right) \] \[ \left( \begin{array}{ccc} 1 & 1 & 1 & 0\\ 0 & 2 & -2 & 0 \\ 0 & 1 & -1 & 0 \end{array} \right) \] \[ \left( \begin{array}{ccc} 1 & 1 & 1 & 0\\ 0 & 1 & -1 & 0 \\ 0 & 1 & -1 & 0 \end{array} \right) \] \[ \left( \begin{array}{ccc} 1 & 1 & 1 & 0\\ 0 & 1 & -1 & 0 \\ 0 & 0 & 0 & 0 \end{array} \right) \] \[ \left( \begin{array}{ccc} 1 & 0 & 2 & 0\\ 0 & 1 & -1 & 0 \\ 0 & 0 & 0 & 0 \end{array} \right) \]

ganeshie8 (ganeshie8):

that gives you x+2z = 0 y-z = 0 z is free

ganeshie8 (ganeshie8):

corresponding eigen vector for \(\lambda = 2\) \[ \left( \begin{array}{ccc} -2z \\ z \\ z \end{array} \right) \]

ganeshie8 (ganeshie8):

\[ \left( \begin{array}{ccc} -2 \\ 1 \\ 1 \end{array} \right) \]

OpenStudy (anonymous):

how did you get this? \left( \begin{array}{ccc} 1 & 0 & 2 & 0\\ 0 & 1 & -1 & 0 \\ 0 & 0 & 0 & 0 \end{array} \right)

OpenStudy (anonymous):

The first row.

ganeshie8 (ganeshie8):

R1-R2

OpenStudy (anonymous):

Okay, understood, but why did you do that?

ganeshie8 (ganeshie8):

i wanted to express x and y in terms of free variable z

ganeshie8 (ganeshie8):

so, in the x row, ive simply eliminated y

OpenStudy (anonymous):

hmm, that makes sense.

ganeshie8 (ganeshie8):

cool/

OpenStudy (anonymous):

How about this? \left( \begin{array}{ccc} -2z \\ z \\ z \end{array} \right)

OpenStudy (anonymous):

Why the z-s?

ganeshie8 (ganeshie8):

cuz our last row is all 0s. z is the free variable here

ganeshie8 (ganeshie8):

you will get infinite multiples for that vector by substituting a value for z

OpenStudy (anonymous):

So, that makes it one since x=2z, right?

ganeshie8 (ganeshie8):

x = -2z dint get ur q ?

OpenStudy (anonymous):

oops, that was an oversight

ganeshie8 (ganeshie8):

hey im not getting u hmm

OpenStudy (anonymous):

Since x = -2z, z=1, right?

ganeshie8 (ganeshie8):

we have the vector in terms of z like below :- -2z z z

ganeshie8 (ganeshie8):

put z = 1, 2, 3, 4.... u wil get different vectors all corresponding to eigen value of 2

ganeshie8 (ganeshie8):

put z =1 -2 1 1 is the eigen vector.

ganeshie8 (ganeshie8):

Also every non-zero multiple of this vector is an eigen vector corresponding to \(\lambda = 2 \)

OpenStudy (anonymous):

Thanks a lot. This really got me crazy for the past week.

ganeshie8 (ganeshie8):

np :) you may use wolframalpha when you're stuck

ganeshie8 (ganeshie8):

here is eigen values and eigen vectors for given matrix from wolfram :- http://www.wolframalpha.com/input/?i=eigen+vectors+of+%7B%282%2C2%2C-2%29+%2C+%281%2C3%2C1%29%2C+%281%2C2%2C2%29%7D

ganeshie8 (ganeshie8):

good luck !

OpenStudy (anonymous):

One more thing

OpenStudy (anonymous):

that question was in terms of b. i just needs clarifications on adjustments to find a or c.

ganeshie8 (ganeshie8):

that im not sure wat they mean by a, b or c i think they mean a, b, c are vectors corresponding to each root

OpenStudy (anonymous):

Oh, I see. Thanks again.

ganeshie8 (ganeshie8):

yw :)